A piggy bank has 100 coins, all of which are nickels and quarters. The total value of the money is $17.80. How many of each denomination are there in the bank?
Answer by TimothyLamb(4379)

x = nickels
y = quarters
---
x + y = 100
5x + 25y = 1780
---
put the system of linear equations into standard form
---
x + y = 100
5x + 25y = 1780
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = nickels = 36
y = quarters = 64
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---